Determine the molality of 3460.00 grams of solution containing 73.3 grams of Potassium Chloride


1
Expert's answer
2022-03-10T10:44:10-0500

Solution:

potassium chloride = KCl

The molar mass of KCl is 74.55 g/mol

Therefore,

Moles of KCl = (73.3 g KCl) × (1 mol KCl / 74.55 g KCl) = 0.9832 mol KCl


Mass of solution = Mass of solvent + Mass of solute

Therefore,

Mass of solvent = Mass of solution − Mass of solute = 3460.00 g − 73.3 g = 3386.7 g


Convert g to kg:

Kilograms of solvent = (3386.7 g) × (1 kg / 1000 g) = 3.3867 kg


Molality = Moles of solute / Kilograms of solvent

Therefore,

Molality of KCl solution = Moles of KCl / Kilograms of solvent

Molality of KCl solution = (0.9832 mol) / (3.3867 kg) = 0.290 mol/kg = 0.290 m

Molality of KCl solution = 0.290 m


Answer: The molality of KCl solution is 0.290 m